# MD-Splatting: Learning Metric Deformation from 4D Gaussians in Highly Deformable Scenes

### v1

@@ 5 Method
MD-Splatting achieves high-quality novel view synthesis using a canonical space of Gaussians and a deformation function to deform them to metric space (Section 5.1). To incentivize learning physically plausible deformations, MD-Splatting introduces several regularization terms (Section 5.2).
Figure 3: MD-Splatting uses three adjacent timesteps at every iteration to enforce physics-inspired regularization terms. All Gaussians are deformed to the metric space using the deformation function $F$, and rasterized to compute the photometric loss and its gradients. The positions of the Gaussians are used to compute the regularization terms based on local rigidity, isometry, and conservation of momentum (Section 5.2).
@@
Canonical Neural Voxel Encoding. As with prior work, MD-Splatting learns a deformation function $F$ from a canonical space. We use a neural-voxel encoding to ensure $F$ has sufficient capacity to capture complex deformations. Prior work [17, 64, 16, 36] showed that neural-voxel encodings improve the speed and accuracy of novel view synthesis in dynamic scenes. We leverage HexPlanes [6, 64] to increase capacity for metric tracking and novel view synthesis simultaneously.
@@
Deformation MLP. The deformation MLP takes in the voxel encoding and uses the encoding to deform the canonical Gaussians into metric space. Figure 2 shows the deformation MLP, which infers position, rotation, and a shadow scalar. We choose this set of outputs to model rigid-body transformations of each Gaussian and changes in illumination. Modeling changes in illumination is critical in the presence of shadows. The RGB color of each Gaussian is multiplied by the shadow scalar $s\in[0,1]$, and the shadow scalar is in the range $[0,1]$ by feeding the output of the MLP through a sigmoid activation function.
Next, we deform the Gaussians, modifying their mean positions $\mu$, rotation $R$, and color $C$, and arrive at a set of Gaussians in the metric space each with state $P$. The differentiable rasterizer from Gaussian Splatting [25] then renders the Gaussians to retrieve gradients for regressing both the canonical Gaussian states, and the parameters of the deformation function.
Unlike 4D-Gaussians [64], we propose to not infer opacity or scale using the deformation field. Optimizing for opacity and scale over time would allow Gaussians to disappear or appear instead of following the motion, which would make tracking less accurate. This design choice reduces the capacity of the deformation function, hence a lower view reconstruction quality as compared to 4D-Gaussians might be expected, see Table 2 for more information.
@@
Motivation. Figure 3 shows the process of tracking Gaussians from the canonical space in metric space. By querying the deformation function $F$, we can track the position of a Gaussian along the entire trajectory.
Without additional supervision, this approach will not necessarily converge to metric deformations. Especially when objects include areas with little texture and uniform color, the solution space for all metric deformations is underconstrained by photometric consistency supervision. To better learn a metric deformation, we propose regularization terms inspired by physics.
After empirically evaluating several combinations of regularization terms, we adopt the rigidity and isometry losses proposed in Luiten et al. and add a momentum term. The first two terms capture a local rigidity loss and local isometry loss, which we compute based on the state of the $k$ nearest neighboring (KNN) Gaussians.
Local Rigidity Loss We assume the deformable material behaves linearly locally. With a large number of Gaussians, dense in 3D space, we expect nearest neighbors to move in similar directions and undergo similar rotations. Without this assumption, the trajectories of Gaussians are free to intersect and cross without a penalty. Instead, we introduce a local rigidity loss,
$\mathcal{L}_{i,j,t}^{\text{rigid}}=w_{i,j}\|(\mu_{j,t-1}-\mu_{i,t-1})-R_{i,t-1}R_{i,t}^{-1}(\mu_{j,t}-\mu_{i,t})\|_{2},$
where $\mu_{i,t}$ is the mean position of Gaussian $i$ at time $t$, $R_{i,t}$ is the rotation matrix of Gaussian $i$ at timestep $t$. This loss regularizes the change in position of the $k$ nearest neighbors as seen from Gaussian $i$, by transforming the current offset of the neighbors to the previous Gaussian orientation. We define
@@
an isotropic Gaussian used to decrease the weight of Gaussians at a greater distance from $\mu_{i,0}$. When a sufficiently large $\lambda_{w}$ is applied, and with a small enough $k$ for KNN, this loss fits the local linearity assumption. The total rigidity loss at time $t$ is
$\mathcal{L}^{\text{rigid}}_{t}=\frac{1}{k|\mathcal{P}|}\sum_{i\in\mathcal{P}}\sum_{j\in\text{knn}_{i}}\mathcal{L}_{i,j,t}^{\text{rigid}}.$
@@
Local Isometry Loss Another assumption we incentivize is to keep the relative position of the $k$ nearest neighbors constant w.r.t. $t=0$. With sufficient deformation, this assumption will be broken at a larger scale, but at a local scale, this regularization avoids drift from the ground truth trajectory. The isometry loss is
Conservation of Momentum The final term we add is to incentivize the conservation of momentum. Newton’s first law states objects without external forces applied, given some mass $m$ and velocity vector $\mathbf{v}$, maintain their momentum $m\cdot\mathbf{v}$. We introduce the regularization term
@@
This term incentivizes a constant velocity vector and has the effect of imposing a low-pass filter on the 3D trajectories. It smooths out trajectories with many sudden changes of direction and magnitude (momentum). Table 1 shows the impact of the proposed regularization terms on tracking errors.
@@ 6 Experiments
We test the effectiveness of MD-Splatting on synthetic datasets of deformable objects. Section 6.1 provides details on our experiment setup, evaluation metrics, and baseline methods. Section 6.2 reports evaluation results from the compared methods and result analysis. Section 6.3 supplies additional ablation results on the training efficiency of MD-Splatting.
6.1 Experiment Setup
Metric Method Scene 1 Scene 2 Scene 3 Scene 4 Scene 5 Scene 6 Mean
3D MTE [mm] $\downarrow$ RAFT [51] ${}^{a}$ 89.687 151.363 295.66 135.468 125.554 67.861 144.266
RAFT Oracle [51]${}^{a}$${}^{b}$ 27.328 34.629 104.842 24.976 32.331 36.113 43.370
OmniMotion Oracle [60] ${}^{a}$${}^{b}$ 16.188 23.51 92.536 11.064 57.603 19.624 36.754
DeVRF [30] 147.293 78.487 449.74 123.745 77.736 133.391 168.399
DynaGS [33] 17.189 54.409 61.148 19.870 24.081 26.757 33.909
4D-GS [64] 7.788 93.409 118.692 16.422 7.033 4.600 41.324
MD-Splatting-NV 5.173 63.894 81.917 9.449 5.679 3.385 28.250
MD-Splatting-TK 2.881 46.257 88.169 9.686 4.741 3.175 25.818
3D $\delta_{\mathrm{avg}}~{}\uparrow$ RAFT [51] ${}^{a}$ 0.442 0.229 0.128 0.295 0.395 0.434 0.321
RAFT Oracle [51] ${}^{a}$${}^{b}$ 0.623 0.562 0.247 0.651 0.573 0.509 0.528
OmniMotion Oracle [60] ${}^{a}$${}^{b}$ 0.763 0.612 0.301 0.763 0.448 0.680 0.595
DeVRF [30] 0.344 0.368 0.094 0.326 0.471 0.398 0.334
DynaGS [33] 0.774 0.401 0.317 0.697 0.649 0.607 0.574
4D-GS [64] 0.787 0.337 0.257 0.700 0.817 0.883 0.630
MD-Splatting-NV 0.851 0.428 0.346 0.749 0.870 0.919 0.694
MD-Splatting-TK 0.912 0.511 0.338 0.780 0.903 0.939 0.731
3D Survival $\uparrow$ RAFT [51] ${}^{a}$ 0.935 0.891 0.738 0.919 0.807 0.891 0.864
RAFT Oracle [51] ${}^{a}$${}^{b}$ 1.000 1.000 1.000 1.000 0.999 0.974 0.996
OmniMotion Oracle [60] ${}^{a}$${}^{b}$ 1.000 0.993 1.000 0.991 0.870 1.000 0.976
DeVRF [30] 1.000 0.972 0.571 0.952 0.860 0.862 0.870
DynaGS [33] 1.000 1.000 0.958 1.000 1.000 1.000 0.993
4D-GS [64] 1.000 0.978 0.967 1.000 0.999 1.000 0.991
MD-Splatting-NV 0.999 1.000 0.987 1.000 1.000 1.000 0.998
MD-Splatting-TK 1.000 1.000 1.000 1.000 1.000 1.000 1.000
${}^{a}$ Method had access to ground truth depth. ${}^{b}$ Method had access to ground truth trajectories to pick the best camera view.
Table 1: 3D tracking results on the deformable cloth dataset (Figure 1). For each metric, the methods above the solid line had access to privileged information, see ${}^{a}$${}^{b}$ and Section 6.1 for more details. The results suggest that MD-Splatting outperforms the baselines in all averaged metrics. DynaGS [33] and OmniMotion [60] are both competitive but accumulate large tracking errors in scenes 1, 5, and 6. The results also show the regularization parameters improve the tracking performance compared to 4D-GS [64].
@@ 6.1 Experiment Setup
Dataset Preparation We use Blender to model dynamic cloth sequences and render photo-realistic images. We create 6 distinct scenes, each containing a different cloth with distinct visual and physical properties, and we render images from 100 different camera views and 40 consecutive time steps for training for a total of 4,000 images. The cloth deformations are introduced by dropping each cloth over one or a few invisible balls onto a ground plane or by constraining the cloth at an attachment point. We obtain ground-truth trajectories by tracking the mesh vertices of deformable objects in Blender.
@@
Baselines. We compare to baseline methods: (1) 2D tracking approaches by projecting their inferences to 3D points using Blender’s ground truth depth and the camera’s ground truth projective geometry. While these methods were not designed for 3D tracking, they are well-known for their impressive 2D tracking performance. Their numbers aid in putting the tracking performance of the other baselines into context. We run RAFT [51] on all views and report the mean results as the RAFT model. We also provide two oracle methods which have access to privileged information: the ground truth trajectories. RAFT Oracle first evaluates on all views, to then output only the trajectories from the view with the lowest median trajectory error. We also report OmniMotion Oracle, which runs OmniMotion [60] on the frame with the lowest MTE for RAFT. Training OmniMotion takes roughly 12–13 hours on an Nvidia RTX 4090 GPU, making inference on all 100 views impractical. The numbers from RAFT Oracle and OmniMotion Oracle are not an apples-to-apples comparison with the other baselines, as to obtain their result they had to access privileged ground-truth trajectories.
(2) DeVRF [30] represents a dynamic, non-rigid scene with a 3D canonical voxel space and a 4D deformation voxel field. It proposes a 4D deformation cycle that models both the backward deformation from the deformed space to the canonical space and the forward deformation from the canonical space to the deformed space. Therefore, the point trajectories can be obtained by querying the forward deformation of canonical points at $t=0$ to every timestep’s deformed space through the 4D forward deformation space. While DeVRF was not designed for 3D tracking directly, it provides useful insight into how similar methods based on NeRF might perform in highly deformable scenes.
@@
(3) Dynamic 3D Gaussians (DynaGS[33], which also builds upon 3D Gaussian splatting for dynamic novel view synthesis, except it explicitly models the positions and rotations of each Gaussian at each time-step. This results in straightforward tracking of any point via finding the trajectory of the learned Gaussian with the highest influence on the given point. Although the original paper assumes a known point cloud at the first frame, we instead use a randomly sampled point cloud for a fair comparison, with DynaGS and MD-Splatting both not using depth information. We also ran with a point cloud initialized using ground-truth depth from Blender but saw very little to no improvement.
@@
(4) Finally, we compare to tracking using 4D-Gaussians [64] (4D-GS). We add our approach for tracking canonical Gaussian, as shown in Figure 3, to extract 3D trajectories from a learning view synthesis model. Comparing to 4D-GS serves to show the impact of the changes made in the model architecture and the regularization terms, to arrive at MD-Splatting.
Metrics Method Scene1 Scene2 Scene3 Scene 4 Scene5 Scene 6 Mean
[…]
DynaGS [33] 35.51 32.23 31.36 30.54 29.84 30.48 31.66
4D-GS [64] 40.76 40.14 44.38 42.05 33.70 34.83 39.31
MD-Splatting-NV 40.67 39.62 43.27 42.61 33.46 34.96 39.10
MD-Splatting-TK 38.76 38.16 40.34 40.61 32.47 32.39 37.12
[…]
DynaGS [33] 0.96 0.93 0.95 0.96 0.92 0.95 0.94
4D-GS [64] 0.99 0.97 0.98 0.99 0.96 0.98 0.98
MD-Splatting-NV 0.99 0.97 0.98 1.00 0.96 0.98 0.98
MD-Splatting-TK 0.99 0.97 0.98 0.99 0.95 0.97 0.98
[…]
DynaGS [33] 0.03 0.11 0.16 0.07 0.10 0.06 0.09
4D-GS [64] 0.01 0.07 0.07 0.02 0.07 0.02 0.04
MD-Splatting-NV 0.01 0.07 0.07 0.02 0.08 0.02 0.05
MD-Splatting-TK 0.02 0.08 0.08 0.03 0.08 0.04 0.05
Table 2: Novel view synthesis results on the deformable cloth dataset (Figure 1). The results suggest MD-Splatting significantly outperforms DynaGS [33] and DeVRF [30], while 4D-GS [64] performs slightly better. This is expected as a result of the additional regularization terms and deformation function design (Section 5). MD-Splatting only takes a minor decrease in view reconstruction performance, in order to significantly improve tracking. The results also suggest that MD-Splatting-TK performs worse in tracking as compared to MD-Splatting-NV, due to the stronger isometry regularization (Section 6.4).
@@
We evaluate 3D tracking and novel-view synthesis results. For 3D tracking we evaluate each compared method on 1,000 randomly sampled points on each cloth. For novel-view synthesis, we evaluate using 200 unseen test-time cameras, and average the results across all cameras and all 40 time steps.
Figure 4: Qualitative novel view synthesis results from all baselines and MD-Splatting. Both 4D-GS [64] and MD-Splatting achieve the best qualitative results that are close to ground-truth renderings (especially on Scene 6, where the floral patterns on the cloth are reconstructed), albeit still fail to capture the fine-grained texture and wrinkles on Scene 2 and Scene3 cloth.
6.2 Main Results
@@
3D Point Tracking. Following prior work [71, 33], we report median trajectory error (MTE), position accuracy ($\delta$), and the survival rate with a threshold of 0.5 [m] [33].
The results are summarized in Table 1. We make the following observations: (1) MD-Splatting outperforms the baselines in all averaged metrics, with a 23.9% better average MTE compared to the baseline methods. Dynamic 3D Gaussians [33], OmniMotion [60], and RAFT [51] are competitive but accumulate a significantly larger tracking error in scenes 1, 4, 5, and 6. For example, in scene 1, the baselines have a MTE of at least 170% higher as compared to MD-Splatting. Excluding our implementation of 4D-GS [64] for tracking, the clostest baseline is the OmniMotion [60] oracle model with a MTE 462% higher as compared to MD-Splatting. (2) MD-Splatting-TK performs better tracking compared to MD-Splatting NV, except for scene 3. The results suggest that setting $\lambda^{\text{iso}}=\text{1.0}$ leads to overall better tracking performance. Section 6.4 contains a more in-depth analysis on this parameter.
(3) MD-Splatting achieves the best performance in environments with textured cloth: scenes 1, 5, and 6. The texture makes photometric consistency a stronger supervision signal, leading to more accurate 3D tracking.
(4) The regularization parameters and change in deformation function architecture seem to improve the tracking performance compared to our implementation of 4D-GS [64]. MD-Splatting scores better metrics in all scenes. (5) The RAFT [51] and OmniMotion [60] results are highly dependent on the camera view angles, and we report the camera view with the best error values after evaluating all the test-time views; despite this given advantage, these methods still fail to achieve consistent and competitive tracking performance.
We hypothesize that part of the tracking error from Dynamic 3D Gaussians [33] is a result of converging Gaussian states to local optima. Especially with larger deformations and shadows, we imagine Gaussians could converge to local optima with relatively large tracking errors. The other source of error we hypothesize plays a role is the various regularization terms and their hyperparameters, which are predominantly focused on rigid motion.
(a) DeVRF [30]
@@
Figure 5: Results on Scene 6: randomly sampled ground-truth trajectories in green, inferred trajectories in red, and the error in red. The differences between 4D-GS [64] are marked with green squares. Compared to the baseline methods, MD-Splatting results in fewer errors in 3D tracking.
Novel View Synthesis. We evaluate visual reconstruction quality. In Table 2, we report results on the three standard metrics commonly used in prior work: PSNR [61], SSIM [62], and LPIPS [70]. MD-Splatting-NV performs comparably to 4D Gaussians (4D-GS) on SSIM while obtaining marginally lower results on PSNR and LPIPS. DynaGS performs worse on view synthesis by a considerable margin. DeVRF obtains the lowest performance, and it does worse on the Scene 2 and 3 as compared to the other scenes—this suggests the textureless cloths impose additional challenges that the model cannot capture. MD-Splatting-TK yields a lower quality view reconstruction as compared to MD-Splatting-NV, as a result of the stronger isometry regularization (Section 6.4).
We jointly analyze results from 3D tracking 1 and novel view sythesis 2. We observe that 4D-GS under-performs DynaGS at 3D tracking in median trajectory error, but outperforms DynaGS at novel view synthesis. MD-Splatting is designed to achieve the best of both worlds. While close to 4D-GS in nature, MD-Splatting can track deformations better than DynaGS while keeping the high-quality novel-view synthesis capability.
Qualitative Results: Tracked Trajectories Figure 5 shows a quantitative comparison for tracking performance in Scene 6, overlaid on each method’ renders. The randomly sampled trajectories show examples of trajectories where DeVRF [30] and Dynamic 3D Gaussians [33] diverge from the ground truth. For DeVRF [30], we hypothesize a lack of high-quality view reconstruction is the cause of poor tracking, whereas Dyamic 3D Gaussians [33] has high-quality view reconstruction but has a few trajectories diverge from the ground truth. The difference between 4D-GS and MD-Splatting-NV is more subtle, and can be observed in the areas marked with green boxes. The green boxes highlight areas where 4D-GS [64] struggles to maintain tracking, whereas MD-Splatting-NV achieves better tracking.
Qualitative Results: Novel View Synthesis. Figure 4 shows a qualitative comparison for scenes 2, 3, and 6. Both 4D-GS and MD-Splatting-NV achieve the best qualitative results that are close to ground-truth (notably on Scene 6, where the floral patterns on the cloth are reconstructed), albeit still fail to capture the fine-grained texture and wrinkles on Scene 2 and 3 cloth. DynaGS also achieves high-quality rendering overall but display a blurry visual effect around the deformed (i.e., wrinkled) areas. We hypothesize that regressing Gaussian positions and orientations is more unstable, and prone to converge to local optima. DeVRF nearly completely fails in Scene 3 due to the lack of textures: the model overfits to one small patch on the cloth and ignores the rest of the scene; it also shows difficulty in learning the large deformations on edges, i.e. only capturing colors around the center of the cloth.
6.3 Training Efficiency Ablation
We compare the training efficiency between MD-Splatting-NV and Dynamic 3D Gaussians [33] when decreasing the number of time steps provided from each camera view. We evaluate on Scene 5 and report PSNR on the subsampled frames, training time, and median trajectory error in Table 3. We observe that for MD-Splatting, training time is less affected by the number of frames from each camera view, as compared to Dynamic 3D Gaussians [33]. This is a result of its approach to train one time step at a time. The results also suggest that MD-Splatting has more robustness to larger deformations between frames, leading to a better PSNR and median trajectory error.
Metric Method 40 Steps 20 Steps 14 Steps 10 Steps

### v2

@@ 5 Method
DeformGS achieves novel-view synthesis and high-quality 3D tracking using a canonical space of Gaussians and a deformation function to deform them to world space (Section 5.1). To incentivize learning physically plausible deformations, DeformGS introduces several regularization terms (Section 5.2). Finally, DeformGS learns 3D masks to focus regularization and Gaussian deformation on dynamic parts of the scene (Section 5.3).
@@
Canonical Neural Voxel Encoding. As with prior work, DeformGS learns a deformation function $F$ from a canonical space. We use a neural-voxel encoding to ensure $F$ has sufficient capacity to capture complex deformations. Prior work [17, 53, 16, 32] showed that neural-voxel encodings improve the speed and accuracy of novel-view synthesis in dynamic scenes. We leverage HexPlanes [6, 53] to increase capacity for simultaneous 3D tracking and novel-view synthesis.
@@
Deformation MLP. The deformation MLP takes in the voxel encoding and uses the encoding to deform the canonical Gaussians into world coordinates. Figure 2 shows the deformation MLP, which infers position, rotation, and a shadow scalar, given a feature vector from the neural voxel encoding. We choose this set of outputs to model rigid-body transformations of each Gaussian and changes in illumination. Modeling changes in illumination is critical in the presence of shadows. We multiply the RGB color of each Gaussian by the shadow scalar $s\in[0,1]$, and the shadow scalar is in the range $[0,1]$ by feeding the output of the MLP through a sigmoid activation function.
Next, we deform the Gaussians, modifying their mean positions $\mu$ and rotation $R$, and arrive at a set of Gaussians in the world space each with state $P$. The differentiable rasterizer from Gaussian Splatting [22] then renders the Gaussians to retrieve gradients for regressing both the canonical Gaussian states and the parameters of the deformation function.
Unlike 4D-Gaussians [53], we propose to not infer opacity or scale using the deformation field. Optimizing for opacity and scale over time would allow Gaussians to disappear or appear instead of following the motion, which would make tracking less accurate. This design choice reduces the capacity of the deformation function, hence a lower view reconstruction quality as compared to 4D-Gaussians might be expected.
@@
Figure 3: DeformGS uses three adjacent timesteps at every iteration to enforce physics-inspired regularization terms. All Gaussians are deformed to world space using the deformation function $F$, and rasterized to compute the photometric loss and its gradients. The positions of the Gaussians are used to compute the regularization terms based on local isometry and conservation of momentum (Section 5.2).
Physics-Inspired Losses. Figure 3 shows the process of tracking Gaussians from the canonical space in world space. By querying the deformation function $F$, we can track the position of a Gaussian along the entire trajectory.
Without additional supervision, this approach will not necessarily converge to physically plausible deformations. Especially when objects include areas with little texture and uniform color, the solution space for all deformations is underconstrained by photometric consistency alone. To learn a more grounded deformation function, we propose regularization terms inspired by physics.
After empirically evaluating several combinations of regularization terms, we adopt the isometry loss proposed in [29] and add a conservation of momentum term. The first term captures a local isometry loss, which we compute based on the state of the $k$ nearest neighboring (KNN) Gaussians.
Local Isometry Loss We incentivize the Gaussians to keep the relative position of the $k$ nearest neighbors constant w.r.t. $t=0$. With sufficient deformation, this assumption will be broken at a larger scale, but at a local scale, this regularization avoids drift from the ground-truth trajectory. The isometry loss is
with
@@
Conservation of Momentum We add a term to incentivize conservation of momentum. Newton’s first law states objects without external forces applied, given some mass $m$ and velocity vector $\mathbf{v}$, maintain their momentum $m\cdot\mathbf{v}$. We introduce the regularization term
@@
This term incentivizes a constant-velocity vector and has the effect of imposing a low-pass filter on the 3D trajectories. It smooths out trajectories with many sudden changes of direction and magnitude (momentum).
5.3 Learning 3D Masks
Learning accurate 3D tracking in scenes with a mix of static and dynamic objects and rich textures poses significant challenges, mainly: (1) imposing physics-inspired regularization terms on all Gaussians may cause issues when dynamic and static objects interact, and (2) modeling millions of dynamic Gaussians can become a significant computational burden.
To address this, DeformGS takes noisy masks of dynamic scene components such as cloth, and learns what Gaussians are dynamic. More formally, we render a mask $M$ by
$M=\sum_{i\in N}m_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$
where $m_{i}$ is a per-Gaussian property, with $m_{i}\in[0,1]$. We then add a regularization term to the loss function s.t. $m_{i}$ is regressed to best reconstruct M. Finally, DeformGS uses $m_{i}$ to select a subset of Gaussians to be dynamic, and applies regularization terms only to those Gaussians.
@@ 6 Experiments
We evaluate DeformGS on synthetic and real-world datasets of scenes with highly deformable objects. Section 6.1 provides details on the simulation experiment setup, evaluation metrics, and baseline methods. Section 6.2 reports evaluation results from the compared methods and provides analysis. Section 6.3 lists the real-world evaluation setup, and finally in Section 6.4 we provide a qualitative evaluation of the performance of DeformGS in the real-world.
6.1 Simulation Experiment Setup
@@
Dataset Preparation We use Blender to model dynamic cloth sequences and render photo-realistic images. We create 6 distinct scenes, each containing a different cloth with distinct visual and physical properties, and we render images from 100 different camera views and 40 consecutive time steps for training for a total of 4,000 images. The cloth deformations are introduced by dropping each cloth over one or a few invisible balls onto a ground plane or by constraining the cloth at an attachment point. We obtain ground-truth trajectories by tracking the mesh vertices of deformable objects in Blender. Every scene contains a single deformable object and a rendered background.
@@
Oracle Baselines We compare DeformGS to 2D tracking oracle models which have access to ground truth depth and trajectory information. While these methods were not designed for 3D tracking, they are well-known for their impressive 2D tracking performance. Their numbers aid in putting the tracking performance of the other baselines into context. We run RAFT [43] on all views, project tracking to 3D using ground truth depth, and report the mean results as the RAFT model. We provide two additional oracle methods which have access to the ground-truth trajectories as well. RAFT Oracle first evaluates on all views, to then output only the trajectories from the view with the lowest median trajectory error. We also report OmniMotion Oracle, which runs OmniMotion [48] on the viewpoint with the lowest MTE for RAFT. Training OmniMotion takes roughly 12–13 hours on an Nvidia RTX 4090 GPU, making inference on all 100 views impractical. The numbers from RAFT Oracle and OmniMotion Oracle are not an apples-to-apples comparison with the other methods, as to obtain their result they had to access privileged ground-truth trajectories.
@@
Gaussian Splatting Baselines (1) Dynamic 3D Gaussians (DynaGS[29], which also builds on 3D Gaussian splatting for dynamic novel-view synthesis, except it explicitly models the positions and rotations of each Gaussian at each time-step. This results in straightforward tracking of any point via finding the trajectory of the learned Gaussian closest to a queried point. Although the original paper assumes a known point cloud at the first frame, we instead use a randomly sampled point cloud for a fair comparison, with DynaGS and DeformGS both not using depth information.
@@
(2) Finally, we compare to tracking using 4D-Gaussians [53] (4D-GS). We add the approach for 3D tracking of canonical Gaussian, as shown in Figure 3, to extract 3D trajectories from a learning view synthesis model. Comparing to 4D-GS serves to show the impact of the changes made in the model architecture, the regularization terms, and using learning per-Gaussian masks to arrive at DeformGS.
@@
We evaluate each compared method on 1,000 randomly sampled points on each cloth.
6.2 Simulation Results
Metric Method Scene 1 Scene 2 Scene 3 Scene 4 Scene 5 Scene 6 Mean
3D MTE [mm] $\downarrow$ RAFT [43] a 67.264 89.944 220.125 177.909 84.593 23.422 110.543
RAFT Oracle [43]ab 3.381 26.956 58.971 12.481 3.930 3.192 18.152
OmniMotion Oracle [48] ab 0.535 14.513 39.958 4.556 2.487 2.011 10.677
DynaGS [29] 24.233 81.119 464.64 54.074 34.985 36.101 115.859
4D-GS [53] 4.645 95.032 223.839 27.441 14.091 11.619 62.778
DeformGS 3.373 45.33 88.369 14.022 7.257 8.173 27.754
3D $\delta_{\mathrm{avg}}~{}\uparrow$ RAFT [43] a 0.553 0.379 0.222 0.533 0.586 0.619 0.482
RAFT Oracle [43]ab 0.926 0.577 0.411 0.703 0.833 0.808 0.710
OmniMotion Oracle [48] ab 0.987 0.693 0.549 0.864 0.885 0.849 0.805
DynaGS [29] 0.624 0.31 0.042 0.435 0.533 0.527 0.4118
4D-GS [53] 0.902 0.339 0.164 0.583 0.697 0.715 0.5667
DeformGS 0.929 0.522 0.322 0.71 0.856 0.787 0.688
3D Survival $\uparrow$ RAFT [43] a 0.945 0.792 0.779 0.822 0.792 0.854 0.831
RAFT Oracle [43]ab 0.986 0.833 0.957 0.872 0.929 0.903 0.913
OmniMotion Oracle [48] ab 1 0.963 1 0.985 0.963 0.933 0.977
DynaGS [29] 0.999 0.99 0.483 0.988 0.992 0.992 0.907
4D-GS [53] 1 0.967 0.834 1 0.999 1 0.967
DeformGS 1 1 1 1 1 1 1
a Method had access to ground truth depth. b Method had access to ground truth trajectories to pick the best camera view.
Table 1: 3D tracking results on the deformable cloth dataset (Figure 1). For each metric, the methods above the solid line had access to privileged information, see ab and Section 6.1 for more details. The results suggest that DeformGS outperforms the baselines in all averaged metrics, and is competitive with the oracle models. The results also suggest our novel deformation function architecture, learning per-Gaussian masks, and physics-inspired regularization losses improve the tracking performance compared to 4D-GS [53]. We do not consider the oracle methods to be fair baselines and therefore do not bold their results.
@@ 6.2 Simulation Results
3D Point Tracking Following prior work [57, 29], we report median trajectory error (MTE), position accuracy ($\delta$), and the survival rate with a threshold of 0.5 [m] [29].
The results are summarized in Table 1. We make the following observations: (1) DeformGS outperforms baselines RAFT, DynaGS, and 4D-GS, by achieving a MTE of 55.8% - 76.0% lower compared to the baselines. (2) The discrepancy between the RAFT oracle model and its averaged result demonstrates the difficulty arising from frequent self-occlusions. This also points to future research avenues for additional supervision through optic flow and 2D tracking algorithms such as RAFT. (3) The oracle models perform very well, this is in part thanks to the falling and short-horizon nature of these sequences, limiting self-occlusions. In the real-world we expect much larger errors due to noisy depth and more challenging occlusions in long-horizon tasks. It would also be unclear what viewpoint to choose without access to ground truth trajectories. (4) Scenes with less texture such as scene 3 perform significantly worse than scenes with strong texture.
Qualitative Results Figure 4 shows ground truth and inferred trajectories in scene 5. The results show that especially DynaGS and 4D-GS introduce large errors as the cloth drapes down. RAFT improves over DynaGS and 4D-GS but requires accurate depth estimation.
(a) RAFT [43]
@@
Figure 4: Results on Scene 5: randomly sampled ground-truth trajectories in green, inferred trajectories in red, and the error of corresponding points in red lines. Compared to the baseline methods, DeformGS results in fewer errors in 3D tracking.
6.3 Real-World Experiment Setup
Robo360 Data The Robo360 dataset [26] is a 3D omnispective multi-material robotic manipulation dataset. It covers many different scenario’s, including manipulation by robot manipulators and humans captured by 86 calibrated cameras. These properties make it an ideal dataset to evaluate the effectiveness of DeformGS in the real world. We select two scenes: (1) a human folding a larger duvet and (2) a human folding a smaller cloth. In the cloth folding scene, we exclude viewpoints where the entire person’s body is visible to eliminate unnecessary complexity.
We also subsample the data to demonstrate DeformGS performance with fewer views. The duvet folding scene contains 17 training views and the cloth folding scene contains 20 training views.
6.4 Real-World Experiment Results
Real2Sim for Digital Twins Figure 5 shows the 3D tracking overlaid on rendered images, as well as the Gaussian points at each time step. The results suggest that DeformGS is able to successfully infer smooth and meaningful trajectories in the real world. While no ground truth is available, the trajectories appear to follow their geometry closely except for a few floating Gaussians. Hyperparameter tuning of the regularization functions, as well as discarding Gaussians with a low opacity, might help resolve this.
The point cloud included in this Figure can be used to create a digital twin after recording the sequence. The digital twin of the duvet, and the entire environment, can then be used to create more dense supervision for imitation learning approaches.
Task-Relevant Keypoint Tracking Robotic manipulators can benefit from tracking task-relevant keypoints, such as the corner of a cloth or the edge of a jacket. Figure 6 shows a comparison between 4D-GS [53] and DeformGS in 3D point tracking, evaluated on both duvet and cloth scenes. The results suggest DeformGS leads to more smooth and overall useful trajectories. The trajectories from 4D-GS intertwine into more messy trajectories, and appear less physically plausible. This would hinder the adoption of 3D tracking into robot applications.
Figure 5: A person manipulating a duvet in the Robo360 [26] dataset, reconstructed using DeformGS. The top row shows the 4D Gaussians as point clouds, where the color represents dense correspondences. The bottom row shows rendered views overlaid with 3D trajectories projected to image space.
Figure 6: Real-world results comparing our proposed DeformGS against 4D-GS [53]. The 3D trajectories inferred by DeformGS appear more smooth and accurate, whereas 4D-GS displays more cluttered trajectories.
7 Conclusions
In this work, we address the challenging problem of 3D point-tracking in dynamic scenes with deformable objects. We introduced DeformGS, the first approach that learns continuous deformations for 3D tracking of deformable scenes. We empirically demonstrate that DeformGS outperforms baseline methods and achieves both high-quality dynamic scene reconstruction and high-accuracy 3D tracking on highly deformed cloth objects with occlusions and shadows, both in simulation and the real world. We also contribute a dataset of six synthetic scenes to facilitate future research.
Limitations and Future Work DeformGS, similar to prior work on dynamic novel view reconstruction, requires a setup of multiple synchronized and calibrated cameras, which may require a significant engineering effort in real-world scenarios. Additionally, significant innovation will be required to achieve the demonstrated results in real-time, as will be beneficial for scalable robot applications.
While DeformGS improves upon prior methods, we do observe Gaussians wandering off in some cases. We also notice the algorithm is relatively sensitive to the regularization hyper parameters ($\lambda^{\text{momentum}}$ and $\lambda^{\text{iso}}$), this might be resolved in the future by adding supervision from state-of-the-art point-tracking algorithms. These limitations point to promising directions for future research.
Acknowledgements This work was supported by the Center for Machine Learning and Health (CMLH) at CMU, and the Pittsburgh Super Computing Center (PSC). We thank David Held for the productive discussions.
{credits}